a rectangles width is one-third its length and its perimeter is 112 m. find the dimensions
Answer by macston(5194)

W=width; L=length=3W; P=2(L+W)=112m
P=2(L+W)
112m=2(L+W) Divide each side by 2.
56m=L+W Substitute for L.
56m=3W+W
56m=4W Divide each side by W.
14m=W ANSWER 1: The width is 14 meters.
L=3W=3(14m)=42m ANSWER 2: The length is 42 meters.
CHECK:
P=2(L+W)
112m=2(42m+14m)
112m=2(56m)
112m=112m